Circular No. 7050

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998es IN NGC 632 E. Halderson, M. Modjaz, T. Shefler, J. Y. King, M. Papenkova, W. D.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova in NGC 632 during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998es was discovered and confirmed on unfiltered observations taken on Nov. 13.3 (mag ∼ 14.6) and 14.4 UT (mag ∼ 14.3). The new object is located at α = 1^{h}37^{m}17^{s}.50, δ = +5^{o}52′50′′.3 (equinox 2000.0), which is ∼ 11′′.4 north of the nucleus of NGC 632. KAIT images of the same field taken on Oct. 24.3 (limiting mag ∼ 16.0) and Oct 21.3 (limiting mag ∼ 19.5) showed nothing at the position of the supernova.

S. Nakano, Sumoto, Japan, reports the following position end figures and magnitude measured by M. Aoki, Toyama, from a CCD image taken on Nov. 14.387 UT: 17^{s}.47, 50′′.3, V = 14.5. Aoki adds that a star of mag 15 is located 10′′ northeast of SN 1998es. Nakano also forwards the following CCD position end figures and unfiltered magnitude for SN 1998es from R. and Y. Kushida, Yatsugatake South Base Observatory (0.40-m f/6.4 Schmidt-Cassegrain reflector) on Nov. 14.510: 17^{s}.48, 50′′.2, 14.5; the offset from the galaxy center is 0′′.4 west and 10′′.8 north. R. Kushida adds that there is no star around the discovery position on her two previous frames of this galaxy taken on 1997 Oct. 10 and Nov. 11.

COMET P/1998 U3 (JÄGER)

Orbital elements from MPC 32866:

T = 1999 Mar. 7.7714 TT  ω = 179.^{o}4942
e = 0.652672  Ω = 303.8178 }2000.0
q = 2.152631 AU  i = 19.0944

a = 6.197683 AU n^{o} = 0.0638793 P = 15.4 years

COMET C/1998 U5 (LINEAR) Visual m_{1} and coma-diameter estimates: Nov. 12.07 UT, 8.7, 7′ (J.

Bortle, Stormville, NY, 20×80 binoculars); 12.67, 8.3, 5′ (T. Seki, Geisei, Japan, 0.20-m refractor); 13.04, 8.2, 9′ (Bortle); 14.07, 7.8, 13′ (Bortle, 10×50 binoculars).

1998 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green